OpenStudy (anonymous):

what is 5 and 4 ninths divided by 7

OpenStudy (jack1):

\(\huge \frac {5\frac 49}7\) yeah??

OpenStudy (anonymous):

what is the anwser if the fraction is 

OpenStudy (jack1):

\(\huge \frac {5\frac 49}7\) \(\huge \frac {\frac {45}9 +\frac 49}7\) \(\huge \frac {\frac {49}9}7\) \(\huge \frac 79\)

OpenStudy (anonymous):

thank you

OpenStudy (jack1):

welcome ;)